OpenStudy (anonymous):

Suppose that you have been hired at an annual salary of $85,000 and receive a raise of 6% per year. How much will you have earned after 7 yrs???

OpenStudy (anonymous):

1 year: 90,100 2 years:95,506 3 years: 101,236.36 So each year, the salary is raised 6%. So changing the percent to a decimal is .06. Using that you would times it with $85,000 and get $5,100. Then you would add that and get the next year's salary. So taking that salary which would be $90,100. Basically, you're multiplying .06 with the previous salary for each year and that is how you would get the next salary. Use this to find out the rest of the years until you reach 7 years.

OpenStudy (tkhunny):

i = 0.06 r = 1 + i = 1.06 \(85000 + 85000r + 85000r^{2} + ... + 85000r^{6} = 85000(1 + r + r^{2} + ... + r^{6}) = 85000\cdot\dfrac{1-r^{7}}{1-r} \) It's called "Basic Principles" and it used the concept of the Sum of a Geometric Series.

OpenStudy (anonymous):

thanks!! I got it now, final answer is $713,476

OpenStudy (anonymous):

Welcome :)

OpenStudy (tkhunny):

That last one was \(85000\cdot\dfrac{1-r^{7}}{1-r}\)